# Generalized and Efficient 2D Gaussian Splatting for Arbitrary-scale Super-Resolution

### v2

@@ Abstract
Equipped with the continuous representation capability of Multi-Layer Perceptron (MLP), Implicit Neural Representation (INR) has been successfully employed for Arbitrary-scale Super-Resolution (ASR). However, the limited receptive field of the linear layers in MLP restricts the representation capability of INR, while it is computationally expensive to query the MLP numerous times to render each pixel. Recently, Gaussian Splatting (GS) has shown its advantages over INR in both visual quality and rendering speed in 3D tasks, which motivates us to explore whether GS can be employed for the ASR task. However, directly applying GS to ASR is exceptionally challenging because the original GS is an optimization-based method through overfitting each single scene, while in ASR we aim to learn a single model that can generalize to different images and scaling factors. We overcome these challenges by developing two novel techniques. Firstly, to generalize GS for ASR, we elaborately design an architecture to predict the corresponding image-conditioned Gaussians of the input low-resolution image in a feed-forward manner. Secondly, we implement an efficient differentiable 2D GPU/CUDA-based scale-aware rasterization to render super-resolved images by sampling discrete RGB values from the predicted contiguous Gaussians. Via end-to-end training, our optimized network, namely GSASR, can perform ASR for any image and unseen scaling factors. Extensive experiments validate the effectiveness of our proposed method. The project page can be found at https://mt-cly.github.io/GSASR.github.io/.
@@ 1 Introduction
Inspired by the development of Implicit Neural Representation (INR) [48], Chen et al. presented LIIF [10] to address the problem of Arbitrary-scale SR (ASR) by making use of the continuity of Multi-Layer Perceptron (MLP) to query the RGB values of a high-resolution (HR) image and feeding its index to the MLP network. The following works try to improve the ASR performance with local feature estimators [32] or scale-aware attention modules [2]. Most of the existing INR-based ASR models [10, 32, 69, 2] rely heavily on MLP layers to predict RGB pixels. However, this implicit function lacks enough representation capacity, as the linear layers in MLP have a very limited receptive field, which leads to limited reconstruction performance. Moreover, since the computations of different indices are usually unshared during the MLP feed-forward process, existing INR-based ASR methods have low computational efficiency and difficulties in scaling up the model parameters.
@@
Recently, Gaussian Splatting (GS) [26] has shown its advantages over INR-based methods, such as Neural Radiance Field (NeRF) [48], in terms of both accuracy and complexity, thanks to its high computational efficiency and powerful explicit representation capabilities. GS has been successfully applied to tasks such as 3D reconstruction [19, 4], representation [49, 18], editing [56, 11], generation [41, 12], etc. While some studies have explored the use of GS for low-level vision tasks, e.g., image deblurring [8, 81], image compression [78], video super-resolution [68], these models overfit on the repetitive images from multi-frames in a video clip or multi-view cameras, or overfit on a single image to complete the image compression task. None of these methods could be applied to single image SR tasks, since overfitting a model on a single image cannot meet the generalization requirements of SR, where a single model should be used to super-resolve any input low-resolution (LR) image to its high-resolution (HR) counterpart.
@@
Instead of overfitting the model to a fixed scenario, we switch from the optimization-based manner of original 3D GS [26] to a learning-based method and present GSASR (Gaussian Splatting based Arbitrary-scale Super-Resolution), which adapts GS to the ASR task. Specifically, we first obtain the deep image features by encoding the input LR image with a general SR backbone [40, 79]. Secondly, inspired by DETR [3], we introduce learnable Gaussian embedding to output 2D Gaussians conditioned on the extracted image features. To enhance interaction among Gaussian embeddings, we perform window self-attention [42] for them so that they can adapt to each other. Then, the Gaussian embeddings are projected to predict the corresponding properties of 2D Gaussians, including the opacity, position, standard deviation, correlation coefficient, and peak RGB values. These 2D Gaussians depict the image in the contiguous Gaussian space. To render 2D Gaussians to obtain the super-resolved image with arbitrary scale, we implement an efficient differentiable GPU/CUDA-based rasterization, which receives the predicted 2D Gaussians with an upsampling scale vector as condition. The scale vector controls the sampling density, and all Gaussians are processed efficiently in parallel.
@@
The advantages of our proposed GSASR are three-folds. (1) Due to the large receptive field brought by window-based self-attention modules, the Gaussian embeddings could sense long-range dependency from deep features and other embeddings. (2) The position of Gaussians are predicted explicitly, enabling Gaussians to concentrate on areas with complex textures, instead of being uniformly distributed over the whole image, thus leading to better reconstruction results. (3) Benefited from the developed efficient scale-aware 2D rasterization technique, GSASR is much faster than the state-of-the-art INR-based models. For example, under the scaling factor ×$\times 12$, GSASR could super-resolve an image to $720\times 720$ resolution within only $91$ ms, while the state-of-the-art INR-based model such as CiaoSR [2] will cost nearly $540$ ms.
@@ 2 Related Work
SR aims to upscale the resolution of an LR input image. The community has witnessed a rapid development of SR methods, ranging from powerful feature extraction backbones [33, 40, 79, 80, 37, 21, 38, 47, 77, 9, 13, 35, 17, 74] to the utilization of generative models [31, 53, 60, 59, 76, 43, 44, 61, 73, 39, 34, 5, 51, 71, 66, 65, 57, 70, 62] for better perceptual quality. While different methods differ from each other on specific designs, they all rely on upsampling operators to magnify the resolution of the original LR image. Various upsampling strategies have been developed, ranging from hand-crafted interpolation operators, to deep neural network-based convolutional layers [16, 54], and to implicit neural representations [10].
@@
Fixed-scale SR Methods. In the early stage, deep neural network-based SR models [15, 27, 28, 30] usually adopt interpolation operators to magnify the resolution. They first upsample the resolution of an LR image to the size of its HR counterpart, and then feed it into network for detail enhancement. Such a scheme is straightforward but suffers from high computational cost. FSRCNN [16] is proposed to apply de-convolutional layers to upsample the resolution of the deep features at the end of the network, so as to accelerate the inference speed. ESPCN [54] performs upsampling by rearranging the LR feature maps into a higher-resolution space through pixel-shuffle layers. However, since the pixels are discrete in feature domain, the de-convolution or pixel-shuffle layers cannot perform continuous sampling to reconstruct SR images with arbitrary scales. Due to these limitations, most of the existing SR methods are conducted on integer scaling factors.
@@
Arbitrary-scale SR Methods. The interpolation operators can be used to continuously predict neighborhood pixels at any magnification factors, yet they will sacrifice the image quality and obtain blurry results. Meta-SR [24] equips the network with the arbitrary-scale magnification ability by designing a meta-upscaling module that leverages DNN to predict HR details from LR neighborhood embeddings. Wang et al. [58] performed arbitrary upsampling by introducing dynamic filters generated based on the scaling factor. Inspired by the success of INR-based methods [48] in 3D tasks, LIIF [10] creatively introduces INR to complete the ASR task by learning an implicit mapping function between a pixel’s RGB value and its index. Following LIIF, many INR-based models have been proposed. ITSRN [67] makes use of a transformer block and implicit position embedding to predict the super-resolved images. LTE [32] employs an implicit function to describe image features in Fourier domain. Based on Neural Operator [20], SRNO [64] adopts a kernel integral operator to fit the implicit function. BTC [50] employs B-spline implicit representation to achieve arbitrary-scale SR of screen content images. LINF [69] estimates a local implicit function with normalizing flow to model the naturalness of textures. CLIT [7] employs a cascaded cross-scale attention block with frequency encoding layers to sense neighborhood features. CiaoSR [2] introduces scale-aware attention-in-attention modules to extend the receptive field to promote the ASR performance. LMF [22] introduces a latent modulated function to compress the feature and accelerate the rendering process at the price of representation accuracy.
@@
Very recently, GaussianSR [23] was proposed to employ GS for the ASR task. However, it has some significant drawbacks. Firstly, it only utilizes the position of RGB pixels from the low-resolution image to substite the center positions of Gaussians, heavily limiting the representation capability of Gaussians and leading to poor reconstruction performance. Secondly, it lacks a differentiable GPU/CUDA-based operator to complete the rasterization process, leading to low-efficiency and high computational cost. In contrast, our proposed GSASR predicts all of the parameters of Gaussians in an end-to-end manner, and we present an efficient differentiable 2D GPU/CUDA-based scare-aware rasterization to largely reduce the computational cost, leading to an effective and efficient model for accomplishing the ASR task.
@@ 3 Method
2D GS Representation. Our key insight is to represent an image using contiguous Gaussians, which can fit various frequency components of an image due to the powerful expressive properties of GS. Specifically, considering an LR image $I_{LR}$ of size $H\times W$, we represent it using $N$ 2D Gaussians, denoted by $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$. Each Gaussian maintains properties of opacity $\alpha~{}\in[0,1]^{1}$, center position $\mu=\{\mu_{x}~{}\in\mathbb{R}^{1},\mu_{y}~{}\in\mathbb{R}^{1}\}$, standard deviation $\sigma=\{\sigma_{x}~{}\in(0,\infty)^{1},\sigma_{y}~{}\in(0,\infty)^{1}\}$, correlation coefficient $\rho~{}\in[-1,1]^{1}$, and normalized peak RGB value $c~{}\in[0,1]^{3}$. We define the color at position $(x,y)$ contributed by Gaussian $G$ as $G(x,y)$ (here we omit the index of Gaussians for simplicity of expression):
@@
$f(x,y)=e^{-\frac{1}{2(1-\rho^{2})}(\frac{\Delta x^{2}}{\sigma_{x}^{2}}-\frac{2\rho\Delta x\Delta y}{\sigma_{x}\sigma_{y}}+\frac{\Delta y^{2}}{\sigma_{y}^{2}})},$ (2)
where $\Delta x=x-\mu_{x}$ and $\Delta y=y-\mu_{y}$. The color of the LR image at position $(x\in[0,W-1],y\in[0,H-1])$ is represented as the summarization of all $N$ Gaussians:
$I_{LR}(x,y)=\sum_{i=1}^{N}G_{i}(x,y).$ (3)
Gaussian functions are contiguous, enabling sampling at arbitrary positions. With a given scaling factor $s$, to obtain an SR image $I_{SR}$ of size $sH\times sW$, we sample the color at position $(x\in[0,sW-1],y\in[0,sH-1])$ with the following scale-aware rasterization:
$I_{SR}(x,y;s)=\sum_{i=1}^{N}G_{i}(\frac{x}{s},\frac{y}{s}).$ (4)
Fast Rasterization. Simply rendering an SR image by querying each pixel from all 2D Gaussians leads to a complexity of $\mathcal{O}(s^{2}HWN)$, which is too high for high-resolution images. Actually, a Gaussian generally focuses on a limited area and its contribution to pixel values decays rapidly with the increase of distance. Therefore, we introduce a rasterization ratio $r<1$ to control the rendering range of each Gaussian. Specifically, we handle all Gaussians in parallel and only render the pixels that are close enough to the Gaussian centers, greatly reducing the computational complexity to $\mathcal{O}(r^{2}s^{2}HWN)$ and making our algorithm practical. ($r$ is set to 0.1 in our implementation.) Note that our rasterization process is differential, which can be seamlessly integrated with neural networks for end-to-end optimization. Algorithm 1 describes the detailed steps of our scale-aware rasterization process. This algorithm is implemented via CUDA C++, which is GPU-friendly and achieves faster speed and low memory requirements.
@@
3
4Initialize $I_{SR}$ as an ($sH$, $sW$, 3) array of zeros.
5For each $G_{i}$ in $\{G_{1},G_{2},\ldots,G_{N}\}$ do
@@ 7       For each pixel $(x,y)$ in $I_{SR}$ do
Table 1: Quantitative comparison between representative ASR models and our GSASR. All models use the same EDSR-backbone [40] as the feature extraction encoder. We test on DIV2K and LSDIR [36] datasets [55] with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$, ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$. The best results are highlighted in red. The PSNR and SSIM indices are computed in the Y channel of Ycbcr space.
@@ 3.2 GSASR: Architecture and Modules
Conventional 3D Gaussian Splatting (GS) methods [26] directly optimize 3D Gaussians to fit a specific scene. As a result, they lack the generalization capability for the ASR tasks, where the model should be generalized to unseen images. To overcome this limitation, we introduce a novel network architecture, which conditions 2D Gaussians on low-resolution (LR) input images, generating content-aware Gaussian representations with enhanced generalization capabilities. Our proposed method, namely GSASR, is illustrated in Figure 2. It begins with an off-the-shelf encoder that extracts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$ from the LR input $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, serving as input to the subsequent Gaussian generation process. The key of GSASR lies in the proposed Gaussian decoder, which receives image feature $\mathbf{F}$ and a learnable Gaussian embedding as inputs, and outputs 2D Gaussians $\mathcal{G}$ to depict the input image in contiguous Gaussian space. Our Gaussian decoder consists of a condition injection block, a Gaussian interaction block, and a Gaussian primary head. Finally, the output of Gaussian decoder is rasterized into the SR image.
Gaussian Embedding. In the GS decoder, we aim to represent the LR image through $N$ Gaussian embeddings $\mathbf{E}\in\mathbb{R}^{N\times d}$ conditioned on $\mathbf{F}$, where $d$ is the embedding dimension. A straightforward idea is to maintain $N$ independent learnable Gaussian embeddings that interact with image features to output $\mathcal{G}$. However, this approach suffers from both effectiveness and computational efficiency. First, the LR images exhibit variable sizes in the test phase. Rendering a fixed number of 2D Gaussians is inadequate to represent the details of a high-resolution LR image. Second, maintaining $N$ independent Gaussian embeddings not only introduces a significant memory burden but also faces optimization challenges, primarily attributed to the initial homogeneity and independent optimization of embeddings.
To address these limitations, we propose two key innovations. First, instead of setting a fixed value for $N$, we let $N$ be proportional to the LR image size ($N\propto H\times W$), employing more Gaussians to represent larger inputs. All Gaussian embeddings share the same initial weights, as shown in Figure 2. Second, to mitigate the optimization complexity and computational demands, inspired by Deform-DETR [82], we assign the $N$ Gaussian embeddings to different reference positions $p\in\mathbb{R}^{N\times 2}$ during initialization, which is obtained by sampling $N$ points at equal intervals from an $H\times W$ image. In this way, different Gaussian embeddings interpret the content at different regions based on their assigned reference positions, while they share the same weight initialization.
@@
Condition Injection Block. As shown in Figure 2, we utilize cross-attention to incorporate feature $\mathbf{F}$ into Gaussian embeddings $\mathbf{E}$. Such a technique is widely used in models like Stable Diffusion [52]. To reduce the computational complexity, we perform window-based cross-attention [42] only in pre-defined local windows as follows:
@@
$\text{Attention}(Q,K,V)=\text{SoftMax}\left(\frac{QK^{\top}}{\sqrt{d}}+B\right)V,$ (5)

### v3

@@ Abstract
Implicit Neural Representation (INR) has been successfully employed for Arbitrary-scale Super-Resolution (ASR). However, INR-based models need to query the multi-layer perceptron module numerous times and render a pixel in each query, resulting in insufficient representation capability and computational efficiency. Recently, Gaussian Splatting (GS) has shown its advantages over INR in both visual quality and rendering speed in 3D tasks, which motivates us to explore whether GS can be employed for the ASR task. However, directly applying GS to ASR is exceptionally challenging because the original GS is an optimization-based method through overfitting each single scene, while in ASR we aim to learn a single model that can generalize to different images and scaling factors. We overcome these challenges by developing two novel techniques. Firstly, to generalize GS for ASR, we elaborately design an architecture to predict the corresponding image-conditioned Gaussians of the input low-resolution image in a feed-forward manner. Each Gaussian can fit the shape and direction of an area of complex textures, showing powerful representation capability. Secondly, we implement an efficient differentiable 2D GPU/CUDA-based scale-aware rasterization to render super-resolved images by sampling discrete RGB values from the predicted continuous Gaussians. Via end-to-end training, our optimized network, namely GSASR, can perform ASR for any image and unseen scaling factors. Extensive experiments validate the effectiveness of our proposed method. The code and models are available at https://github.com/ChrisDud0257/GSASR.
@@ 1 Introduction
Inspired by the development of Implicit Neural Representation (INR) [37], Chen et al. presented LIIF [10] to address the problem of Arbitrary-scale SR (ASR) by making use of the continuity of Multi-Layer Perceptron (MLP) to query the RGB values of a high-resolution (HR) image and feeding its index to the MLP network. The following works try to improve the ASR performance with local feature estimators [27], normalization flow [54] or scale-aware attention modules [2]. Most of the existing INR-based ASR models [10, 27, 54, 2] rely heavily on MLP layers to predict RGB pixels. However, this implicit function lacks enough representation capacity, as it renders the RGB color of a single pixel in each query. INR-based models need to query the MLP layers numerous times, resulting in very low computational efficiency, especially for those models [2] who scale up the MLP layers and employ complex attention modules.
@@
Recently, Gaussian Splatting (GS) [22] has shown its advantages over INR-based methods, such as Neural Radiance Field (NeRF) [37], in terms of both accuracy and complexity, thanks to its high computational efficiency and powerful explicit representation capabilities. GS has been successfully applied to tasks such as 3D reconstruction [17, 4], representation [38, 16], editing [42, 11], generation [33, 12], etc. While some studies have explored the use of GS for low-level vision tasks, e.g., image deblurring [8, 65], image compression [62], video super-resolution [53], these models overfit on the repetitive patterns from multi-frames in a video clip or multi-view cameras, or overfit on a single image to complete the image compression task. None of them could be applied to the image SR task, since overfitting a model on a single image cannot meet the generalization requirements of SR, where a single model should be used to super-resolve any input low-resolution (LR) image to its high-resolution (HR) counterpart.
@@
Instead of overfitting the model to a fixed scenario, we switch from the optimization-based manner of original 3D GS [22] to a learning-based method and present GSASR (Gaussian Splatting based Arbitrary-scale Super-Resolution). Specifically, we first obtain the deep image features by encoding the input LR image with a general SR backbone [32, 63]. Secondly, inspired by DETR [3], we introduce learnable Gaussian embedding to output 2D Gaussians conditioned on the extracted image features. To enhance interaction among Gaussian embeddings, we perform window self-attention [34] so that they can adapt to each other. Then, the Gaussian embeddings are projected to predict the corresponding properties of 2D Gaussians, including the opacity, position, standard deviation, correlation coefficient, and peak RGB values. These 2D Gaussians depict the image in the continuous Gaussian space. To render 2D Gaussians to obtain the super-resolved image with arbitrary scale, we implement an efficient differentiable GPU/CUDA-based rasterization, which receives the predicted 2D Gaussians with an upsampling scale vector as condition. The scale vector controls the sampling density, and all Gaussians are processed efficiently in parallel.
@@
The advantages of our proposed GSASR are three-folds. (1) Since each Gaussian contributes to a large area, GSASR outputs region-level predictions, showing much more powerful representation capability than INR-based models, which only render a single pixel during each query. (2) The attributes of Gaussians are explicitly predicted to ensure better reconstruction results. Specifically, the predicted position enables Gaussians to concentrate on areas with complex textures, while the predicted standard deviation and correlation coefficient indicate the shape and direction of Gaussians, allowing Gaussian to fit the shape of complex textures. (3) Benefited from the developed efficient scale-aware 2D rasterization technique, GSASR is much faster than the state-of-the-art INR-based models. As shown in Fig. 1 and Tab. 2, under the scaling factor ×$\times 12$, GSASR could super-resolve an image to $720\times 720$ resolution within only $91$ ms, while the state-of-the-art INR-based model such as CiaoSR [2] will cost nearly $540$ ms.
@@ 2 Related Work
SR aims to increase the resolution of a given LR image. While different methods differ from each other on specific designs, they all rely on upsampling operators to magnify the resolution of the original LR image, such as hand-crafted interpolation operators, convolutional layers [15, 40], and implicit neural representations [10].
@@
Fixed-scale SR Methods. In the early stage, deep neural network-based SR models [14, 23, 24, 26] usually adopt interpolation operators to magnify the resolution. They first upsample the resolution of an LR image to the size of its HR counterpart, and then feed it into networks for detail enhancement. Such a scheme is straightforward but suffers from high computational cost. FSRCNN [15] is proposed to apply de-convolutional layers to upsample the resolution of the deep features at the end of the network, so as to accelerate the inference speed. ESPCN [40] performs upsampling by rearranging the LR feature maps into a higher-resolution space through pixel-shuffle layers. However, since the pixels are discrete in feature domain, the de-convolution or pixel-shuffle layers cannot perform continuous sampling to reconstruct SR images with arbitrary scales.
@@
Arbitrary-scale SR Methods. The interpolation operators can be used to continuously predict neighborhood pixels at any magnification factors, yet they will sacrifice the image quality and obtain blurry results. Meta-SR [20] designsing a meta-upscaling module that leverages DNN to predict HR details at arbitrary-scales from LR neighborhood embeddings. Wang et al. [44] performed arbitrary upsampling by introducing dynamic filters generated based on the scaling factor. Inspired by the success of INR-based methods [37] in 3D tasks, LIIF [10] introduces INR to complete the ASR task by learning an implicit mapping function between a pixel’s RGB value and its index. ITSRN [52] makes use of a transformer block and implicit position embedding to predict the super-resolved images. LTE [27] employs an implicit function to describe image features in Fourier domain. SRNO [49] adopts a kernel integral operator to fit the implicit function. LINF [54] estimates a local implicit function with normalizing flow to model the naturalness of textures. CLIT [7] employs a cascaded cross-scale attention block with frequency encoding layers to sense neighborhood features. CiaoSR [2] introduces scale-aware attention-in-attention modules to extend the receptive field to promote the ASR performance. LMF [18] introduces a latent modulated function to compress the feature and accelerate the rendering process at the price of representation accuracy.
@@
Very recently, GaussianSR [19] was proposed to employ GS for the ASR task. However, it uses only utilizes the pixel positions of LR image to substitute the center positions of Gaussians, heavily limiting the representation capability of Gaussians. In contrast, our proposed GSASR predicts all of the parameters of Gaussians, and we present an efficient differentiable 2D GPU/CUDA-based scare-aware rasterization to largely reduce the computational cost, leading to an effective and efficient model for accomplishing the ASR task.
@@ 3 Method
2D GS Representation. Our key insight is to represent an image using contiguous Gaussians, which can fit various frequency components of an image due to the powerful expressive properties of GS. Specifically, considering an LR image $I_{LR}$ of size $H\times W$, we represent it using $N$ 2D Gaussians, denoted by $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$. Each Gaussian maintains properties of opacity $\alpha~{}\in[0,1]^{1}$, center position $\mu=\{\mu_{x}~{}\in\mathbb{R}^{1},\mu_{y}~{}\in\mathbb{R}^{1}\}$, standard deviation $\sigma=\{\sigma_{x}~{}\in(0,\infty)^{1},\sigma_{y}~{}\in(0,\infty)^{1}\}$, correlation coefficient $\rho~{}\in[-1,1]^{1}$, and normalized peak RGB value $c~{}\in[0,1]^{3}$. We define the RGB color at position $(x,y)$ contributed by Gaussian $G$ as $G(x,y)$ (here we omit the index of Gaussians for simplicity of expression):
@@
$\vspace{-0.8em}\begin{split}f(x,y)=&{(2\pi\sigma_{x}\sigma_{y}\sqrt{1-\rho^{2}})^{-1}}exp[{-\frac{1}{2(1-\rho^{2})}}\times\\ &{(\frac{\Delta x^{2}}{\sigma_{x}^{2}}-\frac{2\rho\Delta x\Delta y}{\sigma_{x}\sigma_{y}}+\frac{\Delta y^{2}}{\sigma_{y}^{2}})}]\end{split}$ (2)
where $\Delta x=x-\mu_{x}$ and $\Delta y=y-\mu_{y}$ decide the position of Gaussians on an image, and the standard deviation $\sigma_{x}$, $\sigma_{y}$ and the correlation coefficient $\rho$ indicate the shape and direction of Gaussian. The pixel color of the LR image at position $(x\in[0,W-1],y\in[0,H-1])$ is represented as the summarization of all $N$ Gaussians:
$I_{LR}(x,y)=\sum_{i=1}^{N}G_{i}(x,y).\vspace{-0.6em}$ (3)
@@
2
3Initialize $I_{SR}$ as an ($sH$, $sW$, 3) array of zeros.
4For each $G_{i}$ in $\{G_{1},G_{2},\ldots,G_{N}\}$ do
@@ 6       For each pixel $(x,y)$ in $I_{SR}$ do
Gaussian functions $G$ are infinite and continuous on the 2D space, enabling discretely sampling at arbitrary positions. With a given scaling factor $s$, to obtain an SR image $I_{SR}$ of size $sH\times sW$, we sample the color at position $(x\in[0,sW-1],y\in[0,sH-1])$ with the following scale-aware rasterization:
$I_{SR}(x,y;s)=\sum_{i=1}^{N}G_{i}(\frac{x}{s},\frac{y}{s}).\vspace{-0.8em}$ (4)
Fast Rasterization. In Eq. 4, simply rendering an SR image by querying each pixel from all 2D Gaussians leads to a complexity of $\mathcal{O}(s^{2}HWN)$, which is too high for high-resolution images. Actually, a Gaussian generally focuses on a local area and its contribution to pixel values decays rapidly with the increase of distance. Therefore, we introduce a rasterization ratio $r\leq 1$ to control the rendering range of each Gaussian. Specifically, we handle all Gaussians in parallel and only render the pixels that are close enough to the Gaussian centers, greatly reducing the computational complexity to $\mathcal{O}(r^{2}s^{2}HWN)$ and making our algorithm efficient. To achieve a better balance between performance and efficiency, we set the ratio as $r=0.1$, which could significantly speed up the rasterization process and avoid missing major Gaussian responses. Note that our rasterization process is differential, which can be seamlessly integrated into neural network training for end-to-end optimization. Algorithm 1 describes the detailed steps of our scale-aware rasterization process. This algorithm is implemented via CUDA C++, which is GPU-friendly and achieves fast speed and low memory requirements.
Table 1: Quantitative comparison between representative ASR models and our GSASR. All models use the same EDSR-backbone [32] as the feature extraction encoder and are tested on DIV2K and LSDIR [28] datasets [41] with scaling factors ×$\times 2$, ×$\times 3$, ×$\times 4$, ×$\times 6$, ×$\times 8$, ×$\times 12$, ×$\times 16$, ×$\times 18$, ×$\times 24$, ×$\times 30$. The best results are highlighted in red. The PSNR and SSIM indices are computed in the Y channel of Ycbcr space.
@@ 3.2 GSASR: Architecture and Modules
Conventional 3D GS methods [22] directly optimize 3D Gaussians to fit a specific scene. As a result, they lack the generalization capability for the ASR tasks, where the model should be generalized to unseen images. To overcome this limitation, we introduce a novel network architecture, which conditions 2D Gaussians on LR input images, generating content-aware Gaussian representations with enhanced generalization capabilities. Our proposed method, namely GSASR, is illustrated in Figure. 2. It begins with an off-the-shelf encoder (such as EDSR-Baseline [32], RDN [64]) that extracts image features $\mathbf{F}\in\mathbb{R}^{H\times W\times C}$ from the LR input $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$. The deep features $\mathbf{F}$ are input to the subsequent Gaussian generation process. The key of GSASR lies in the proposed Gaussian decoder, which receives image feature $\mathbf{F}$ and learnable Gaussian embeddings as inputs, and outputs 2D Gaussians $\mathcal{G}$ to depict the input image in continuous Gaussian space. Our Gaussian decoder consists of a condition injection block, a Gaussian interaction block, and a Gaussian primary head. Finally, the Gaussian decoder output is rasterized into an SR image.
@@
Gaussian Embedding. In the GS decoder, we aim to represent the LR image through $N$ Gaussian embeddings $\mathbf{E}\in\mathbb{R}^{N\times d}$ conditioned on $\mathbf{F}$, where $d$ is the embedding dimension. As for the number of Gaussian embeddings $N$, a straightforward idea is to set it to a fixed number so that the $N$ independent learnable Gaussian embeddings interact with image features to output 2D Gaussians $\mathcal{G}$. However, this approach suffers from low effectiveness and computational efficiency. As the LR images exhibit variable sizes in the test phase, rendering a fixed number of 2D Gaussians will either waste computational resources when handling LR image with very small size, or lead to deficient representation when processing larger LR images. Another possible strategy is to set $N\propto s$ under different scaling factord $s$. However, it requires the Gaussian decoder to handle varying sizes of Gaussian embeddings, while the size of LR features is fixed in the training stage. In addition, as the scaling factor increases, more Gaussians are introduced, leading to longer inference time and higher GPU memory cost.
To address these issues, we set $N=m\times(H\times W)$ so that it is proportional to the LR image size, where $m$ is a hyper-parameter controlling the density of Gaussians. In this way, we could employ more Gaussians to represent larger inputs, while still keep efficiency under ultra-high scaling factors. To be more specific, we split LR feature $\mathbf{F}\in\mathbb{R}^{C\times H\times W}$ into $\frac{H}{k}\times\frac{W}{k}$ windows with window size $k\times k$. The content in each window is fitted by a learnable Gaussian embedding $\mathbf{E}_{base}\in\mathbb{R}^{mk^{2}\times d}$, which is randomly initialized and optimized during training. To dynamically support arbitrary LR feature size $H\times W$, we duplicate $\mathbf{E}_{base}$ for $\frac{H}{k}\times\frac{W}{k}$ times to obtain $\mathbf{E}\in\mathbb{R}^{N\times d}$ and cover all windows, then we assign the $N$ Gaussian embeddings to different reference positions $p\in\mathbb{R}^{N\times 2}$, which are obtained by sampling $N$ points at equal intervals from an $H\times W$ image. According to reference positions, each $\mathbf{E}_{base}$ will interact with LR features in the window, as shown in Fig. 2.
@@
Condition Injection Block. In this block, each Gaussian embedding $\mathbf{E}_{i},i\in[1,N]$ aims to learn associated features. As shown in Fig. 2, we perform window-based cross-attention [34] to incorporate feature $\mathbf{F}$ into Gaussian embeddings $\mathbf{E}$. Such a technique is widely used in Stable Diffusion [39]. The formulation is as follows:
@@
$\text{Attention}(Q,K,V)=\text{SoftMax}\left(\frac{QK^{\top}}{\sqrt{d}}+B\right)V,\vspace{-0.5em}$ (5)